# Streaming algorithms for products of probabilities

### v1

@@ 3 Lower bound for the product approximation problem
Theorem 3.1 ([1]).
For every $20<\epsilon<1/$ 2 there is a deterministic streaming algorithm for $\mathsf{APP}_{\epsilon}$ with space complexity $\log(n^{2}b/\epsilon)=2\log n+\log b-\log\epsilon$.
Let us briefly explain the idea since it will be needed later. Assume that the input stream is $q_{1}q_{2}\cdots q_{n}$. Define $\epsilon^{\prime}=\epsilon/n$. For an $a\in\mathbb{N}$ we define the bucket $B_{a,\epsilon^{\prime}}=((1-\epsilon^{\prime})^{a+1},(1-\epsilon^{\prime})^{a}]$. Then, the algorithm computes for every input number $q_{i}$ the unique $a_{i}\in\mathbb{N}$ such that $q_{i}\in B_{a_{i},\epsilon^{\prime}}$. Then the streaming algorithmen for $\mathsf{APP}_{\epsilon}$ computes the sum $a=\sum_{i=1}^{n}a_{i}$ and at the end returns $(1-\epsilon^{\prime})^{a}$. It can be shown that this is a $(1-\epsilon)$-approximation of the product of the $q_{i}$. Moreover, every bucket index $a_{i}$ can be bounded by $nb/\epsilon$ (in [1] the bound $4nb/\epsilon$ is stated, but it easy to see that $nb/\epsilon$ suffices.) Hence, every partial sum of the $a_{i}$ is bounded by $n^{2}b/\epsilon$ for which $2\log n+\log b-\log\epsilon$ bits suffice.
The space bound stated in [1] also contains the space needed for computing the memory state transition function (called $\delta_{n,b}$ in Section 2.1). This is mainly the space for computing the bucket indices $a_{i}$. Recall from Section 2.1 that we ignore this internal space in our model. We match the upper bound $2\log n+\log b-\log\epsilon$ by a lower bound up to a factor of two:
Theorem 3.2.
Let $20<\epsilon<1/$ 2. Every deterministic streaming algorithm for $\mathsf{APP}_{\epsilon}$ has space complexity at least $\log n+\log b-\log\epsilon-\Theta(1)$ whenever $-\log\epsilon\leq b<-\log(1-\epsilon)\cdot n/2$.
Proof 3.3.
Fix an $\epsilon$ with $20<\epsilon<1/$ 2 and let $\delta=-\log(1-\epsilon)$. Note that $10<\delta<$ 1. Let $\mathcal{D}$ be a deterministic streaming algorithm for $\mathsf{APP}_{\epsilon}$. It suffices to show that $\mathsf{space}_{\mathcal{D}}(2n,b)\geq\log n+\log b-\log\delta-\Theta(1)$ under the assumption
$-\log\epsilon\leq b<\delta\cdot n.$ (2)
Clearly, we can then subsitute $n$ by $n/2$. Moreover, it can be easily shown that the difference $\log\epsilon-\log\delta=\log\epsilon-\log\log(1/(1-\epsilon))$ belongs to the interval $(-1,\log\ln 2)\subseteq(-1,-0.52)$ when $20<\epsilon<1/$ 2. Thus, we have $\log\epsilon-\Theta(1)=\log\delta-\Theta(1)$.
Due to (2) (which yields 1/$21/2^{b}\leq\epsilon<1/$ 2) there is an integer $k$ with $3\leq k\leq 2^{b}$ and
@@
$S=\bigg{\{}\frac{k-1}{k},\frac{k-2}{k-1},2^{-b}\bigg{\}}.$ (5)
@@
$y:=\left\lfloor\frac{nb}{\delta}\right\rfloor.$
@@
$m=\left\lfloor\frac{j\delta}{b}\right\rfloor\leq\frac{j\delta}{b}\leq\frac{y\delta}{b}\leq n.$
@@
$=$ $(1-\epsilon)^{-n}(1-\epsilon)^{j}\ =\ (1-\epsilon)^{j-n}.$
@@
In the second step we choose the numbers $q_{m+1},\ldots,q_{m+j-s}$ inductively such that $q_{1}q_{2}\cdots q_{m+j-s}\in B_{j}$. Note that $m+j-s\leq 2n$. Assume that $q_{m+1},\ldots,q_{m+d}\in S$ have been chosen for some $0\leq d<j-s$ such that $q_{1}q_{2}\cdots\cdots q_{m+d}\in B_{s+d}$. Note that for $d=0$ this holds by (6). We choose $q_{m+d+1}$ by a case distinction depending on $q_{1}q_{2}\cdots\cdots q_{m+d}$. It might actually happen that $q_{1}q_{2}\cdots\cdots q_{m+d}$ satisfies both cases 1 and 2 below. Then we can choose $q_{m+d+1}$ according to one of them. Recall also that $k$ is such that $3\leq k\leq 2^{b}$ and (4) holds.
@@
$\subseteq$ $\left((1-\epsilon)^{s+d+2},(1-\epsilon)^{s+d+1}\right]=B_{s+d+1}$
where the inclusion follows from $(k-1)/k\geq(1-\epsilon)$; see (4). We can therefore choose $q_{m+d+1}=(k-1)/k$.
@@
where the inclusion in $B_{s+d}$ follows from $(k-1)(1-\epsilon)/(k-2)\geq 1$; see again (4). Also note that $k\geq 3$. Then we get
@@
where the inclusion follows again from (4). We can therefore choose $q_{m+d+1}=(k-2)/(k-1)$.
@@
But this is true, because by (3) and $k\geq 3$ we have $\epsilon\geq k^{-1}\geq(k-1)^{-2}$. This concludes the proof of Claim 1. Consider the buckets $B_{3j}$ for $0\leq j\leq\lfloor y/3\rfloor$. By Claim 1, for every $0\leq j\leq\lfloor y/3\rfloor$ there is a number $a_{j}\in B_{3j}$ such that $a_{j}$ can be written as a product of length at most $2n$ over the set $S$. Let $s_{j}$ be the corresponding input stream whose product is $a_{j}$.
@@
This means that the deterministic streaming algorithm $\mathcal{D}$ for $\mathsf{APP}_{\epsilon}$ must yield different output numbers for the input streams $s_{i}$ and $s_{j}$. In particular, the streaming algorithm must arrive in different memory states for the input streams $s_{i}$ and $s_{j}$. Since this holds for all $0\leq i<j\leq\lfloor y/3\rfloor$, the algorithm must have at least $\lfloor y/3\rfloor+1\geq y/3$ memory states and therefore must store at least $\log(y/3)=\log n+\log b-\log\delta-\Theta(1)$ bits.
@@
Theorem 3.4.
Let $20<\epsilon<1/$ 2. Every randomized streaming algorithm for $\mathsf{APP}_{\epsilon}$ has space complexity $\Omega(\log n+\log b-\log\epsilon-\Theta(1))$ whenever $-\log\epsilon\leq b<-\log(1-\epsilon)\cdot n/2$.
Proof 3.5.
We can show this result by a standard reduction to the one-way communication complexity of $\mathsf{GT}_{m}$; see Theorem 2.1. Let $\mathcal{R}$ be a randomized streaming algorithm for $\mathsf{APP}_{\epsilon}$. As in the proof of Theorem 3.2, we define $\delta=-\log(1-\epsilon)$. It suffices to show that $\mathsf{space}_{\mathcal{R}}(2n,b)\geq\Omega(\log n+\log b-\log\delta-\Theta(1))$ under the assumption (2). We can assume that the error probability of $\mathcal{R}$ is bounded by $1/5$. Consider the number $y$ and the buckets $B_{5i}$ for $0\leq i\leq\lfloor y/5\rfloor$ from the proof of Theorem 3.2. Let $s_{i}$ be an input stream whose product yields a number $a_{i}\in B_{5i}$.
We show that $\mathcal{R}$ yields a randomized one-way protocol for $\mathsf{GT}_{\lfloor y/5\rfloor}$. Assume that $i$ is the input for Alice and $j$ is the input for Bob, where $0\leq i,j\leq\lfloor y/5\rfloor$. Alice starts by running the streaming algorithm $\mathcal{R}$ on the input stream $s_{i}$. Thereby she uses here random bits to simulate the probabilistic choices of $\mathcal{R}$. Let $\alpha_{i}$ by the final memory state of $\mathcal{R}$ obtained by Alice. She sends $\alpha_{i}$ to Bob. Bob then runs $\mathcal{R}$ on input $s_{j}$; let $\alpha_{j}$ be the final memory state of $\mathcal{R}$ obtained by Bob. Bob then applies the output function of $\mathcal{R}$ to the memory states $\alpha_{i}$ and $\alpha_{j}$ and obtains rational numbers $r_{i}$ and $r_{j}$ respectively. With probability at least $(5/6)^{2}\geq 2/3$ the output numbers $r_{i}$ and $r_{j}$ satisfy $a_{i}(1-\epsilon)\leq r_{i}\leq a_{i}/(1-\epsilon)$ and $a_{j}(1-\epsilon)\leq r_{j}\leq a_{j}/(1-\epsilon)$, and thus $r_{i}\in B_{5i+1}\cup B_{5i}\cup B_{5i-1}$ and $r_{j}\in B_{5j+1}\cup B_{5j}\cup B_{5j-1}$. This allows Bob to distinguish the cases $i<j$ and $i\geq j$:
@@
This shows that the communicated memory state $\alpha_{i}$ must have bit length $\Omega(\log(y/5))=\Omega(\log n+\log b-\log\delta-\Theta(1))$.
@@ 4 The threshold problem
Theorem 4.1.
Let $B=2^{b}$. There are deterministic streaming algorithms for $\mathsf{TPP}$ with space complexity $2nb$, respectively $\mathcal{O}(B)$.
Proof 4.2.
For notational convenience we consider streams of length $n+1$. Consider such an input stream $q_{0}q_{1}\cdots q_{n}\in\mathbb{P}^{*}$ of $n+1$ rational probabilities with $\|q_{i}\|\leq b$ for all $0\leq i\leq n$. Recall that $q_{0}$ is the threshold value. A streaming algorithm can simply store all the $q_{i}$ with at most $2(n+1)b$ bits in total.
Another way to store the product $\prod_{i=1}^{k}q_{i}$ for $1\leq k\leq n$ is to store for every prime $p\leq B$ how often it appears in the product, where an occurrence in the denominator is counted negative. There are $\mathcal{O}(B/\log B)=\mathcal{O}(B/b)$ primes $p\leq B$ and the total number of occurrences of a prime $p$ in the product is between $-nb$ and $nb$. This leads to the space bound $\mathcal{O}(B(\log n+\log b)/b)$.
This can be improved to $\mathcal{O}(B)$ as follows: We can assume that $q_{0}>0$, otherwise one only has to check whether one of the $q_{i}$ ($1\leq i\leq n$) is zero. Since $\|q_{0}\|\leq b$ we must have $q_{0}\geq 1/B$. Moreover, we can ignore all $1$’s in the stream $q_{1}q_{2}\cdots q_{n}$. If $q_{i}<1$ then $q_{i}\leq(B-1)/B$. Assume now that the number of $i$ with $1\leq i\leq n$ and $q_{i}\leq(B-1)/B$ is at least $B\ln B$. Then, the product $\prod_{i=1}^{n}q_{i}$ is upper bounded by
$\left(\frac{B-1}{B}\right)^{B\ln B}<\left(\frac{1}{e}\right)^{\ln B}=\frac{1}{B}\leq q_{0}.$
Hence, the algorithm can safely output $1$.
This means that we can replace in the above space bound $\mathcal{O}(B(\log n+\log b)/b)$ the value $n$ by $B\ln B\leq\mathcal{O}(Bb)$. We obtain the space bound $\mathcal{O}(B)$ (this also includes the space for storing $q_{0}$).
Theorem 4.3.
Fix a (small) constant $\gamma>0$. Every deterministic streaming algorithm $\mathcal{D}$ for $\mathsf{TPP}$ satisfies $\mathsf{space}_{\mathcal{D}}(2n,(2+\gamma)(\log_{2}n+b))\geq nb$ for $n$ and $b$ large enough.
Proof 4.4.
Let $B=2^{b}$. We take the first $nB+n+1$ prime numbers $p_{1},p_{2},...,p_{nB+n+1}$. It is known that $p_{k}\leq k\cdot(\ln k+\ln\ln k)$ for $k\geq 6$ [7, 3.13]. In all numbers appearing in the two input streams that we will construct, the numerator and denominator are bounded by $p_{nB+n+1}^{2}$. Therefore, the bit size of all the numbers involved is bounded by
$\lceil\log_{2}p_{nB+n+1}^{2}\rceil=\lceil 2\log_{2}p_{nB+n+1}\rceil\leq(2+\gamma)(\log_{2}n+b)$
if $n$ and $b$ are large enough (depending on the constant $\gamma$).
For every $1\leq i\leq n$ we define the set
@@
Thus, $Q_{i}$ consists of $B$ fractions of two consecutive primes. We then consider the set of all words
$S_{n,b}=\{q_{1}q_{2}\cdots q_{n}\in\mathbb{P}^{n}:q_{i}\in Q_{i}\text{ for all}1\leq i\leq n\}.$
Clearly, $|S_{n,b}|=B^{n}$. We set the threshold (the first number in the stream) to
$t=\frac{p_{1}}{p_{nB+n+1}}.$
Assume now that there is a deterministic streaming algorithm $\mathcal{D}$ for $\mathsf{TPP}$ such that $\mathsf{space}_{\mathcal{D}}(2n,(2+\gamma)(\log_{2}n+b))<n\cdot b=\log_{2}|S_{n,b}|$ bits for $n$ and $b$ large enough. Then there must exist two different words $t\,q_{1}q_{2}\cdots q_{n},\,t\,r_{1}r_{2}\cdots r_{n}\in S_{n,b}$ such that after reading these two words, $\mathcal{D}$ must be in the same memory state. W.l.o.g. we assume that $\prod_{i=1}^{n}q_{i}<\prod_{i=1}^{n}r_{i}$.
We continue the words $tq_{1}q_{2}\cdots q_{n}$ and $tr_{1}r_{2}\cdots r_{n}$, respectively, with the length-$n$ word $(s_{1}/r_{1})(s_{2}/r_{2})\cdots(s_{n}/r_{n})$, where
$s_{i}=\frac{p_{(i-1)(B+1)+1}}{p_{i(B+1)+1}}.$
@@
$\prod_{i=1}^{n}s_{i}=\prod_{i=1}^{n}\frac{p_{(i-1)(B+1)+1}}{p_{i(B+1)+1}}=t.$
@@
But this is a contradiction, since $\mathcal{D}$ arrives in the same memory state after reading the two words.
Theorem 4.5.

### v2

@@ 3 Lower bound for the product approximation problem
Theorem 3.1([2])
For every $0<\epsilon<1/2$ there is a deterministic streaming algorithm for $\mathsf{APP}_{\epsilon}$ with space complexity $\log(n^{2}b/\epsilon)=2\log n+\log b-\log\epsilon$.
Let us briefly explain the idea since it will be needed in Section 5. Assume that the input stream is $q_{1}q_{2}\cdots q_{n}$. Define $\epsilon^{\prime}=\epsilon/n$. For an $a\in\mathbb{N}$ we define the interval $B_{a,\epsilon^{\prime}}=((1-\epsilon^{\prime})^{a+1},(1-\epsilon^{\prime})^{a}]$, called a bucket in [2]. The streaming algorithm computes for every input number $q_{i}$ the unique $a_{i}\in\mathbb{N}$ such that $q_{i}\in B_{a_{i},\epsilon^{\prime}}$ and computes the sum $a=\sum_{i=1}^{n}a_{i}$. At the end it returns $(1-\epsilon^{\prime})^{a}$. It can be shown that this is a $(1-\epsilon)$-approximation of the product of the $q_{i}$. Moreover, every bucket index $a_{i}$ can be bounded by $nb/\epsilon$ (in [2] the bound $4nb/\epsilon$ is stated, but it easy to see that $nb/\epsilon$ suffices.) Hence, every partial sum of the $a_{i}$ is bounded by $n^{2}b/\epsilon$ for which $2\log n+\log b-\log\epsilon$ bits suffice.
The space bound stated in [2] also contains the space needed for computing the memory state transition function (called $\delta_{n,b}$ in Section 2.1). This is mainly the space for computing the bucket indices $a_{i}$. Recall from Section 2.1 that we ignore this internal space in our model. We match the upper bound $2\log n+\log b-\log\epsilon$ by a lower bound up to a factor of two for a reasonable range of the parameters $n$, $b$, and $\epsilon$.
Theorem 3.2
Let $0<\epsilon<1/2$. Every deterministic streaming algorithm for $\mathsf{APP}_{\epsilon}$ has space complexity at least $\log n+\log b-\log\epsilon-\Theta(1)$ when $n$, $b$, and $\epsilon$ satisfy $\log\epsilon\leq b<-\log(1-\epsilon)\cdot n/2$.
Proof
Fix an $\epsilon$ with $0<\epsilon<1/2$ and let $\delta=-\log(1-\epsilon)$. Note that $0<\delta<1$. Let $\mathcal{D}$ be a deterministic streaming algorithm for $\mathsf{APP}_{\epsilon}$. It suffices to show that $\mathsf{space}_{\mathcal{D}}(2n,b)\geq\log n+\log b-\log\delta-\Theta(1)$ under the assumption
$\log\epsilon\leq b<\delta\cdot n.$ (3)
We can then subsitute $n$ by $n/2$. Moreover, it can be easily shown that the difference $\log\epsilon-\log\delta=\log\epsilon-\log\log(1/(1-\epsilon))$ belongs to the interval $(-1,\log\ln 2)\subseteq(-1,-0.52)$ when $0<\epsilon<1/2$. Thus, we have $\log\epsilon-\Theta(1)=\log\delta-\Theta(1)$.
By (3) (which yields $1/2^{b}\leq\epsilon<1/2$) there is a $k\in\mathbb{N}$ with $3\leq k\leq 2^{b}$ and
@@
$S=\bigg\{\frac{k-1}{k},\frac{k-2}{k-1},\frac{1}{2^{b}}\bigg\}\subseteq\mathbb{P}.$ (6)
@@
$y:=\left\lfloor\frac{nb}{\delta}\right\rfloor.$ (7)
@@
[…]
@@
$\ =\$$(1-\epsilon)^{-n}(1-\epsilon)^{j}\ =\ (1-\epsilon)^{j-n}.$
@@
In the second step we choose the numbers $q_{m+1},\ldots,q_{m+j-s}$ inductively such that $q_{1}q_{2}\cdots q_{m+j-s}\in B_{j}$. Note that $m+j-s\leq 2n$. Assume that $q_{m+1},\ldots,q_{m+d}\in S$ have been chosen for some $0\leq d<j-s$ such that $q_{1}q_{2}\cdots\cdots q_{m+d}\in B_{s+d}$. Note that for $d=0$ this holds by (8). We choose $q_{m+d+1}$ by a case distinction depending on $q_{1}q_{2}\cdots\cdots q_{m+d}$. It might happen that $q_{1}q_{2}\cdots\cdots q_{m+d}$ satisfies both cases 1 and 2 below. Then we can choose $q_{m+d+1}$ according to one of them. Recall also that $k$ is such that $3\leq k\leq 2^{b}$ and (5) holds.
@@
where the inclusion in the second line follows from $(k-1)/k\geq(1-\epsilon)$; see (5). We can therefore choose $q_{m+d+1}=(k-1)/k\in S$.
@@
where the inclusion in $B_{s+d}$ follows from $(k-1)(1-\epsilon)/(k-2)\geq 1$; see again (5). Also note that $k\geq 3$ so that k2k-2 is not zero. Then we get
@@
where the inclusion in the second line follows again from (5). We can therefore choose $q_{m+d+1}=(k-2)/(k-1)\in S$.
@@
But this is true, because by (4) and $k\geq 3$ we have $\epsilon\geq k^{-1}\geq(k-1)^{-2}$. This concludes the proof of the claim. ∎
Consider the buckets $B_{3j}$ for $0\leq j\leq\lfloor y/3\rfloor$. By our claim, for every $0\leq j\leq\lfloor y/3\rfloor$ there is a number $a_{j}\in B_{3j}$ such that $a_{j}$ can be written as a product of length at most $2n$ over the set $S$. Let $s_{j}$ be the corresponding input stream whose product is $a_{j}$.
@@
This means that the deterministic streaming algorithm $\mathcal{D}$ for $\mathsf{APP}_{\epsilon}$ must yield different output numbers for the input streams $s_{i}$ and $s_{j}$. In particular, the streaming algorithm must arrive in different memory states for the input streams $s_{i}$ and $s_{j}$. Since this holds for all $0\leq i<j\leq\lfloor y/3\rfloor$, the algorithm must have at least $\lfloor y/3\rfloor+1\geq y/3$ memory states and therefore must store at least $\log(y/3)=\log n+\log b-\log\delta-\Theta(1)$ bits. ∎
@@
Theorem 3.3
Let $0<\epsilon<1/2$. Every randomized streaming algorithm for $\mathsf{APP}_{\epsilon}$ has space complexity $\Omega(\log n+\log b-\log\epsilon-\Theta(1))$ when $n$, $b$, and $\epsilon$ satisfy $\log\epsilon\leq b<-\log(1-\epsilon)\cdot n/2$.
Proof
We can show this result by a standard reduction to the one-way communication complexity of $\mathsf{GT}_{m}$; see Theorem 2.1. Let $\mathcal{R}$ be a randomized streaming algorithm for $\mathsf{APP}_{\epsilon}$. As in the proof of Theorem 3.2, we define $\delta=-\log(1-\epsilon)$. It suffices to show that $\mathsf{space}_{\mathcal{R}}(2n,b)\geq\Omega(\log n+\log b-\log\delta-\Theta(1))$ under the assumption (3). We can assume that the error probability of $\mathcal{R}$ is bounded by $1/6$. Consider the number $y$ and the buckets $B_{5i}$ for $0\leq i\leq\lfloor y/5\rfloor$ from the proof of Theorem 3.2. Let $s_{i}$ be an input stream of length at most $2n$ whose product yields a number $a_{i}\in B_{5i}$.
We show that $\mathcal{R}$ yields a randomized one-way protocol for $\mathsf{GT}_{\lfloor y/5\rfloor+1}$. Assume that $i$ is the input for Alice and $j$ is the input for Bob, where $0\leq i,j\leq\lfloor y/5\rfloor$. Alice starts by running the streaming algorithm $\mathcal{R}$ on the input stream $s_{i}$. Thereby she uses here random bits to simulate the probabilistic choices of $\mathcal{R}$. Let $\alpha_{i}$ by the final memory state of $\mathcal{R}$ obtained by Alice. She sends $\alpha_{i}$ to Bob. Bob then runs $\mathcal{R}$ on input $s_{j}$; let $\alpha_{j}$ be the final memory state of $\mathcal{R}$ obtained by Bob. Bob then applies the output function of $\mathcal{R}$ to the memory states $\alpha_{i}$ and $\alpha_{j}$ and obtains rational numbers $r_{i}$ and $r_{j}$ respectively. With probability at least $(5/6)^{2}\geq 2/3$ the output numbers $r_{i}$ and $r_{j}$ satisfy $a_{i}(1-\epsilon)\leq r_{i}\leq a_{i}/(1-\epsilon)$ and $a_{j}(1-\epsilon)\leq r_{j}\leq a_{j}/(1-\epsilon)$, and thus $r_{i}\in B_{5i+1}\cup B_{5i}\cup B_{5i-1}$ and $r_{j}\in B_{5j+1}\cup B_{5j}\cup B_{5j-1}$. Bob can then distinguish the cases $i<j$, $i\geq j$, and $i>j$:
@@
With Theorem 2.1, this shows that the communicated memory state $\alpha_{i}$ must have bit length $\Omega(\log(y/5))=\Omega(\log n+\log b-\log\delta-\Theta(1))$. ∎
@@ 4 The threshold problem
Theorem 4.1
There are deterministic streaming algorithms for $\mathsf{TPP}$ with space complexity $2nb$ and $1.443\cdot 2^{b}$ (the latter for $b$ large enough), respectively.
Proof
Let $B=2^{b}$. For notational convenience we consider streams of length $n+1$. Consider such an input stream $q_{0}q_{1}\cdots q_{n}\in\mathbb{P}^{*}$ of $n+1$ rational probabilities with $\|q_{i}\|\leq b$ for all $0\leq i\leq n$. Recall that $q_{0}$ is the threshold value. A streaming algorithm for $\mathsf{TPP}$ can simply store all the $q_{i}$ with at most $2(n+1)b$ bits in total.
Another way to store the product $\prod_{i=1}^{k}q_{i}$ for $1\leq k\leq n$ is to store for every prime $p\leq B$ how often it appears in the product, where an occurrence in the denominator is counted negative. By the prime number theorem, the number of primes $p\leq B$ is asymptotically $B/\ln(B)$ and therefore bounded by $1.4427\cdot B/b$ for $b$ large enough (note that $1.4427>\log_{2}(e)$). The total number of occurrences of a prime $p$ in the product is between nb-nb and $nb$. This leads to the space bound $1.4427\cdot B\cdot(\log n+\log b+1)/b$.
This can be improved to $1.443\cdot B$ as follows: We can assume that $q_{0}>0$, otherwise the algorithm can output 0. Since $\|q_{0}\|\leq b$ we must have $q_{0}\geq 1/B$. Moreover, we can ignore all $1$’s in the stream $q_{1}q_{2}\cdots q_{n}$. If $q_{i}<1$ then $q_{i}\leq(B-1)/B$. Assume now that the number of $i$ with $1\leq i\leq n$ and $q_{i}\leq(B-1)/B$ is at least $B\cdot\ln B$. Then, the product $\prod_{i=1}^{n}q_{i}$ is upper bounded by
$\left(\frac{B-1}{B}\right)^{B\cdot\ln B}<\left(\frac{1}{e}\right)^{\ln B}=\frac{1}{B}\leq q_{0}.$
Hence, the algorithm can safely output $1$. This means that we can replace in the above space bound $1.4427\cdot B\cdot(\log n+\log b+1)/b$ the value $\log n$ by $\log(B\cdot\ln B)\leq b+\log(b)+\mathcal{O}(1)$ which gives the space bound $1.443\cdot B$ for $b$ large enough (this also includes additional space $\mathcal{O}(b)$ for storing $q_{0}$ and a binary counter up to $B\cdot\ln B$ for the number of $q_{i}\leq(B-1)/B$). ∎
Let us now come to lower bounds for $\mathsf{TPP}$.
Theorem 4.2
Fix a constant $\gamma>0$. Every deterministic streaming algorithm $\mathcal{D}$ for $\mathsf{TPP}$ satisfies $\mathsf{space}_{\mathcal{D}}(2n,(2+\gamma)(\log_{2}n+b))\geq nb$ for $n$ and $b$ large enough depending on $\gamma$.
Proof
Let $B=2^{b}$. We take the first $nB+n+1$ prime numbers $p_{1},p_{2},...,p_{nB+n+1}$. It is known that $p_{k}\leq k\cdot(\ln k+\ln\ln k)$ for $k\geq 6$ [10, 3.13]. In all numbers appearing in the two input streams that we will construct, the numerator and denominator are bounded by $p_{nB+n+1}^{2}$. Therefore, the bit size of all the numbers involved is bounded by
$\lceil\log_{2}p_{nB+n+1}^{2}\rceil=\lceil 2\log_{2}p_{nB+n+1}\rceil\leq(2+\gamma)(\log_{2}n+b)$ (9)
if $n$ and $b$ are large enough (depending on the constant $\gamma$). Assume for the following that (9) holds. For every $1\leq i\leq n$ we define the set
@@
Thus, $Q_{i}$ consists of $B$ fractions of two consecutive primes. Note that the sets $Q_{i}$ are pairwise disjoint. We then consider the set of all words
$S_{n,b}=\{q_{1}q_{2}\cdots q_{n}\in\mathbb{P}^{n}:q_{i}\in Q_{i}\text{ for all }1\leq i\leq n\}.$ (10)
Clearly, $|S_{n,b}|=B^{n}$. Moreover, for two different words $q_{1}q_{2}\cdots q_{n},r_{1}r_{2}\cdots r_{n}\in S_{n,b}$ the products $\prod_{i=1}^{n}q_{i}$ and $\prod_{i=1}^{n}r_{i}$ are also different due to the uniqueness of prime factorizations. We set the threshold (the first number in the stream) to
$t=\frac{p_{1}}{p_{nB+n+1}}.$ (11)
Assume now that there is a deterministic streaming algorithm $\mathcal{D}$ for $\mathsf{TPP}$ such that $\mathsf{space}_{\mathcal{D}}(2n,(2+\gamma)(\log_{2}n+b))<n\cdot b=\log_{2}|S_{n,b}|$. Then there must exist two different words $u=q_{1}q_{2}\cdots q_{n}$ and $v=r_{1}r_{2}\cdots r_{n}$ in $S_{n,b}$ such that after reading $tu$ and $tv$, $\mathcal{D}$ is in the same memory state. W.l.o.g. we assume that $\prod_{i=1}^{n}q_{i}<\prod_{i=1}^{n}r_{i}$. We continue the words $tu$ and $tv$, respectively, with the length-$n$ word $w=(s_{1}/r_{1})(s_{2}/r_{2})\cdots(s_{n}/r_{n})$, where
$s_{i}=\frac{p_{(i-1)(B+1)+1}}{p_{i(B+1)+1}}.$ (12)
@@
$\prod_{i=1}^{n}s_{i}=\prod_{i=1}^{n}\frac{p_{(i-1)(B+1)+1}}{p_{i(B+1)+1}}=\frac{p_{1}}{p_{nB+n+1}}=t.$
@@
But this is a contradiction, since $\mathcal{D}$ arrives in the same memory state after reading the streams tuwtuw and tvwtvw. ∎
From the construction in the proof of Theorem 4.2 we can easily obtain a lower bound for randomized streaming algorithms for $\mathsf{TPP}$:
Theorem 4.3